a car travels from stop a to stop b at speed of 30 km per hour and returns back at speed of 50 kmper hour. find the total distance

Grade:9

1 Answers

Arun
6 years ago
Average speed =total distance /time
let the distance travelled from A to B= x km
then,
total distance travelled = 2x km
time while going from A to B = D/S
ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย =x/30 hours
time while returning = x/50 hours
total time = x/30 + x/50
ย  ย  ย  ย  ย  ย  ย  ย  =5x+ 3xย 
ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  150
ย ย  ย  ย  ย  ย  ย  ย  ย  =8x/150
ย  ย  ย  ย  ย  ย  ย  ย  ย  =4x/75 hours
average speed = 2x *ย 75
ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย 4x
ย ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย 37.5 km/hr